OpenStudy (anonymous):

2r^3+18r^2=20r then I rewrote it as: 2r^3+18r^2-20r=0 now what do I do?

OpenStudy (bahrom7893):

factor out 2r

OpenStudy (anonymous):

r^2(2r+18-20)

OpenStudy (anonymous):

I forgot the =0

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

hello anyone

OpenStudy (anonymous):

wat r u trying to find out here?

OpenStudy (anonymous):

I'm factoring equations

jimthompson5910 (jim_thompson5910):

2r^3+18r^2=20r 2r^3+18r^2-20r = 0 2r(r^2+9r-10) = 0 Now find two numbers that multiply to -10 and add to 9 at the same time

OpenStudy (anonymous):

-1,10

jimthompson5910 (jim_thompson5910):

so r^2+9r-10 factors to (r-1)(r+10) which means 2r(r^2+9r-10) = 0 becomes 2r(r-1)(r+10) = 0

jimthompson5910 (jim_thompson5910):

See what to do from here?

OpenStudy (anonymous):

r-1=0 and r+10=0

OpenStudy (anonymous):

so r=1 and r=-10

jimthompson5910 (jim_thompson5910):

good, but you're forgetting about the 2r out front 2r = 0 2r/2 = 0/2 r = 0

jimthompson5910 (jim_thompson5910):

so r = 0, r = 1, or r = -10

OpenStudy (anonymous):

thank you so very much

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

:)